What score is 1/2 standard deviation above the mean?

A standardized test was given to a set of high school juniors and the distribution of the data is bell shaped. The mean score is 800 and the standard deviation is 120.

Answer by stanbon(75887)


---------------------
A standardized test was given to a set of high school juniors and the distribution of the data is bell shaped. The mean score is 800 and the standard deviation is 120.
----
What score is 1/2 standard deviation above the mean?
z = 1/2
score = 800 + (1/2)120 = 860
-------------
Cheers,
Stan H.
------------